OpenStudy (starryknight):

What is the solution to the following system of equations? free metal for best answer -2x - 4y = 28 -4x + 0y = -8 please help me i dont understand how to do this euqtion

OpenStudy (anonymous):

-2x - 4y = 28 -4x = -8 ( you can remove the 0y since that would be 0) _multiply second row by .5_ -2x - 4y = 28 -2x = -4 Solving for x: x = 2 -4 - 4y = 28 (if -2x = -2) -4y = 32 y = -8 Checking work: -2(2) - 4(-8) = 28 -2(2) - (-32) = 28 -4 + 32 = 28 28 = 28